# FRIEDA: Benchmarking Multi-Step Cartographic Reasoning in Vision-Language Models

### v1

@@ 1 Introduction
Recent advances in large vision-language models (LVLMs) have markedly improved multimodal reasoning, with strong results across diverse visual question-answering (VQA) tasks (c:dong2024insightv; c:souibgui2025docvxqa). Education and cognitive science research characterizes reasoning as a broad capability that spans numeric reasoning, logical deduction (c:holyoak2012thinkingreasoning), and textual interpretation (c:wharton1991textreason), as well as interpreting pictures (c:mayer2020multimedia), spatial data (c:li2025mapqa), and map images (c:goodchild2006esri). Extensive LVLM benchmarks cover many of these facets: visual numeracy in chart and infographics (c:lin2025infochartqa; c:matthew2022infovqa; c:mary2022chartqa), document and layout reasoning (c:duan2025docopilot; c:matthew2021docvqa), multi-image inference (c:kazemi2025remi; c:xia2025mmie), and even spatial relations in natural images (c:shiri2024spatialreasoningqa). However, reasoning over maps, also a core human competence (c:tversky2003navigating; c:kastens2006spatialthinking; c:ishikawa2021spatialspecial), which we refer to as cartographic reasoning, remains under-examined in LVLMs.
@@
Unlike natural images, maps encode information with an abstract, symbolic visual grammar (e.g., map scales, compass/north arrows, and thematic symbology) (c:buckley2012esri), which demands a deeper interpretation than simple pattern recognition. Mastery of these elements must be coupled with the comprehension of spatial relations that are commonly grouped into topological reasoning (e.g., detecting shared boundaries), metric inference (e.g., converting map lengths to real-world distances through the map scale), and directional reasoning (using a compass or north arrow) (c:clementini19939deim; c:harmelen2007spatialrelation). In addition, human map-reading competencies (c:liben2010competencies; c:muir1985competencies) frequently require these inferences across multiple maps within a single document. Correctly answering a map question, therefore, draws on map-specific skills (c:hegarty2005spatialabilities) such as interpreting map elements, reasoning over spatial relations, and integrating evidence across multiple maps, as well as broader capabilities emphasized in LVLM research that include text grounding (c:singh2019textvqa; c:sidorov2019textcaps), numeric and logical inference (c:lu2024mathvista; c:hu2023treeofmixedthought), multi-image integration (c:wang2024muirbench; c:xia2025mmie), and retrieval (c:wang2025mmniah; c:wang2024niamh). A cartographic reasoning benchmark can therefore probe comprehensive reasoning and provide a clear understanding of the spatial intelligence of LVLMs.
@@
A growing line of work began to evaluate LVLMs on map VQA, yet these benchmarks do not fully assess cartographic reasoning. Earlier datasets pose chart-style questions that can be answered without interpreting spatial relations, which bypasses the topological, metric, and directional inferences that are central to map comprehension (c:koukouraki2025mapvqa; c:chang2022mapqa). Other efforts cover only a subset of relations as they target specific tasks such as navigation (c:feng2025canmllmguidemehome; c:kazemi2025remi) or entity identification (c:dihan2025mapeval). While suitable for those objectives, such coverage is insufficient for evaluating human-like map understanding (c:liben2010competencies). Many benchmarks also restrict the stylistic variability of maps. Some focuses on choropleths (c:koukouraki2025mapvqa; c:chang2022mapqa; c:mukhopadhyay2025mapwise), others rely on maps created with map-coloring tools (c:srivastava2025mapiq) or common web basemaps (c:kazemi2025remi; c:dihan2025mapeval). Several further focus on limited thematic domains (e.g., geology (c:huang2025peace)) or restricted geography coverage (c:chang2022mapqa; c:srivastava2025mapiq). These constraints overlook the heterogeneity in toponyms, labeling conventions, projections, and symbology that real-world cartography demands (c:slocum2022thematiccartography; c:robinson1978elementofcartography). Multi-map reasoning is rarely evaluated, with limited exceptions (c:kazemi2025remi), even though practical cases often require integrating evidence across multiple maps (e.g., reconciling transit maps with future land-use maps for urban planning) and aligning overlapping information (c:lupien1987general). Moreover, although document-level multimodal understanding is emphasized in other LVLM benchmarks, existing map VQA benchmarks seldom require selecting the correct map among many images in long reports, despite government documents and technical documents containing numerous, visually similar, context-dependent maps (c:fema; c:eis; c:ni43101). As a result, current map VQA settings underestimate the demands of comprehensive map understanding, leaving it unclear whether LVLMs possess human-like map-reading competencies. Full cartographic reasoning remains beyond the scope of what existing VQA benchmarks assess.
@@
We introduce FRIEDA, a benchmark designed for evaluating multi-map, multi-step, comprehensive cartographic reasoning in LVLMs. We curate maps from public documents of various thematic domains (e.g., geological surveys, planning reports, environmental studies) to develop questions that require models to interpret maps as they appear in reports, mirroring practical scenarios in which a reader must synthesize evidence from maps embedded in a document (see Figure 1). The collection encompasses a diverse range of styles, projections, and scales. We create each question such that it requires (1) reasoning over topological, metric, and directional relations, (2) interpreting map elements and their semantics (e.g., legends, map scales, and north arrows), and, when applicable, (3) integrating information across multiple maps, and (4) selecting the appropriate map(s) from a document to answer the query. To probe genuine reasoning rather than random guessing, the answers are in a free-form (not multiple-choice) format. The benchmark evaluation includes two settings: a direct setting, which provides the relevant map images with the question to focus on evaluating map comprehension, and a contextual setting, where the model must first retrieve the correct maps from a broader within-document collection before answering. We score outputs using a unified, task-aware protocol aligned to the three spatial-relation categories. We evaluate textual responses (topological and semantic labels) with LLM-as-Judge (c:gu2025surveyllmasajudge), distance responses (numeric values with units) with unit-aware parsing and mean absolute percentage error (MAPE), and directional responses (cardinal directions for relative position) with angular tolerance over the eight directions. We compare the result against the human upper bound derived from multi-annotator agreement to contextualize LVLM performance. By aligning our tasks with the competencies expected of human map-readers (c:goodchild2006esri; c:liben2010competencies) and explicitly targeting compositional cross-image inference that is largely absent from prior map VQA work, FRIEDA fills a crucial gap in state-of-the-art LVLM evaluation.
@@ 2 Task Definition
Spatial relations describe how geographic features relate in space (c:carlson2001relation), how they are positioned in space (c:majic2021spatial), and how their geometries interact (c:guo1998relation). In geographic information systems (GIS) and spatial cognition, these relations are often grouped into three categories: topological, metric, and directional (c:harmelen2007spatialrelation; c:clementini19939deim). To make these abilities measurable and comparable, FRIEDA separates questions by spatial relation type and grounds the topological portion in the 9-intersection model (c:clementini19939deim). We consolidate finer-grained subtypes into their broader categories (e.g., cross classified as intersect, and contain classified as within), yielding four topological classes: border (shared boundary between regions), equal (coincident geometries), intersect (crossing or overlap of features), and within (containment or inclusion of one area inside another). We complement these with one metric primitive, distance, and one directional primitive, orientation. Together, these six relations maintain the expressiveness of spatial queries while aligning with users’ intuitive spatial reasoning.
@@
Maps are symbolic representations that encode spatial information through abstract conventions (c:slocum2022thematiccartography). Therefore, interpreting map elements is a distinct skill central to cartographic reasoning. The key elements we target are map text (place and feature names), legends (mappings from color, icons, and patterns to semantic classes), map scales (measurements that convert the map distance to the real-world distance), and the compass (c:arcgis). The styles of these components vary widely across maps: map texts may use different typography or placement rules (c:monmonier2015history), legends may use continuous color ramps or discrete pictograms (c:slocum2022thematiccartography), map scales may appear as bars or frames around the map (c:robinson1995elements), and the compass may be a compass rose or a north arrow (c:slocum2022thematiccartography). Practical map interpretation requires grasping the concepts of map elements rather than simply recognizing their shapes. Consequently, our design includes questions that require reading map texts, decoding legends, using the map scale, and applying orientation to demonstrate true map literacy by linking abstract visual encodings to their underlying semantics.
@@
Beyond interpreting spatial relations and map elements, practitioners regularly perform cross-map comparison and fusion to synthesize multiple map editions or thematic layers (c:lupien1987general). Our multi-map setting reflects this practice: we curate questions that present two or more maps together and require the model to integrate evidence by aligning shared symbols, reconciling differences in labels, map scales, and orientation, and identifying co-referent regions or features (c:foody2007mapcomparison). Extracting distributions and patterns is widely recognized as a core capability (c:ishikawa2016spatialthinking; c:rexigel2024morebetter; c:morita2025multimap). By testing this setting, we move beyond isolated spatial computation to evaluate deeper cartographic reasoning across varied depictions of the same space.
@@
To mirror practical workflows (c:matthew2021docvqa; c:tanaka2023slidevqa), we evaluate a contextual setting (FRIEDA-contextual), where a model must identify the relevant map before answering a question. In this scenario, we provide the model with multiple maps from the same source (i.e., a document), and the model must perform within-document retrieval using cues in the map, such as titles, legends, or labels. By evaluating model performance on FRIEDA-contextual, we capture a core aspect of real map use: the model must understand how map elements encode meaning and leverage that understanding to select the required map from thematically related alternatives that vary in data layers, geographic extent, or purpose (c:ishikawa2016spatialthinking).
@@ 3.1 Benchmark Statistics
Table 1 shows that FRIEDA comprises 17,030 map images drawn from 210 documents. To capture real-world variability, these maps span diverse geographies (32 countries) and six thematic domains, exhibiting heterogeneous styles, including varied color palettes, legends, and symbol conventions. The benchmark contains a total of 500 questions, comprising 202 single-map and 298 multi-map questions. The multi-map subset consists primarily of two-map comparisons (295 questions), with a few cases requiring reasoning across three (2 questions) or four (1 question) maps.
@@
For FRIEDA-contextual, we provide between 2 and 9 irrelevant maps in addition to the relevant map(s) required to answer the question. The total input image averages 9.54 maps ($\sigma=1.27$) per question across both the single- and multi-map settings, with an average relevant-to-irrelevant maps ratio of 1:5.71.
@@ 3.2 Benchmark Construction
To capture stylistic and geographic diversity, we curate maps from publicly available government and multilateral reports across domains including geology (c:ni43101), national park management (c:national-parks), environmental assessments (c:eis; c:ireland; c:seychelles), disaster response (c:fema), urban planning (c:seattle-planning; c:abu-dhabi; c:sg-ura; c:capetown), and infrastructure investment (c:aiib). We limit the sources to documents using the Latin characters to focus on cartographic reasoning over translation. We extract images using Idefics3-8B (c:laurencon2024idefics) with a custom prompt (Appendix B.1) and manually verify that each extracted set contains only cartographic maps (examples of excluded non-maps appear in Appendix B.1.1). To support FRIEDA-contextual, we retain only documents with at least four maps. We select contextual maps based on their page proximity to the target map; this ensures they are thematically and stylistically related to the target maps. We then shuffle the map order to prevent LVLM from using positional cues to identify the target maps.
@@
For each collected map, we use GPT-4 and GPT-o3 (Appendix B.6) to propose candidate questions, the targeted spatial relations, and a reference answer. We do not impose templates and accept any phrasing that unambiguously represents the target relation as valid to reflect various forms of paraphrases of spatial relations (e.g., “Is A within B?” vs. “Does B contain A”). To ensure the benchmark tests visual cartographic reasoning rather than search or memorization, we discard questions answerable by text-only web search or without visual inspection of the map image.
@@
All LLM-proposed candidate questions undergo a pre-annotation curation stage. The two question curators (one with 7 years of GIS experience and another with 2 years of experience in geospatial data) manually verify gold answers against source maps and rewrite or discard ambiguous questions. This step ensures FRIEDA consists only of high-quality, unambiguous questions before the broader annotator validation phase.
@@
We validate every question with annotations from 11 Ph.D. researchers (eight with map expertise) collected over four weeks. Annotators confirm that each question is answerable from the provided map(s) and, for multi-map questions, verify that all maps are required to answer the question. To prevent bias, curators do not validate their own edits. We only keep the question if a majority ($\geq$2/3) agrees with the gold answer. In a rare case (currently two questions in FRIEDA) where all three annotators agree on an answer contradicting the gold answer, we conduct a secondary review to update the gold answer if consensus is reached. In total, we remove 61 questions that fail to reach an agreement $\geq$2/3. Appendix B.3 details the instruction prompt provided to the annotators, and Appendix B.5 shows the annotation interface.

### v2

# arXiv:2512.08016v2[cs.CV] 27 Feb 2026

@@ 1 Introduction
Recent advances in large vision-language models (LVLMs) have markedly improved multimodal reasoning, with strong results across diverse visual question-answering (VQA) tasks (Dong et al., 2024; Souibgui et al., 2025). Education and cognitive science research characterizes reasoning as a broad capability that spans numeric reasoning, logical deduction (Holyoak and Morrison, 2012), and textual interpretation (Wharton and Kintsch, 1991), as well as interpreting pictures (Mayer, 2020), spatial data (Li et al., 2025), and map images (Goodchild, 2012). Extensive LVLM benchmarks cover many of these facets: visual numeracy in chart and infographics (Lin et al., 2025; Mathew et al., 2022; Masry et al., 2022), document and layout reasoning (Duan et al., 2025; Mathew et al., 2021), multi-image inference (Kazemi et al., 2025; Xia et al., 2025), and even spatial relations in natural images (Shiri et al., 2024). However, reasoning over maps, also a core human competence (Tversky, 2003; Kastens and Ishikawa, 2006; Ishikawa and Newcombe, 2021), which we refer to as cartographic reasoning, remains under-examined in LVLMs.
@@
Unlike natural images, maps encode information with an abstract, symbolic visual grammar (e.g., map scales, compass/north arrows, and thematic symbology) (Buckley, 2006), which demands a deeper interpretation than simple pattern recognition. Mastery of these elements must be coupled with the comprehension of spatial relations that are commonly grouped into topological reasoning (e.g., detecting shared boundaries), metric inference (e.g., converting map lengths to real-world distances through the map scale), and directional reasoning (using a compass or north arrow) (Clementini et al., 1993; Cohn and Renz, 2007). In addition, human map-reading competencies (Liben et al., 2010; Muir, 1985) frequently require these inferences across multiple maps within a single document. Correctly answering a map question, therefore, draws on map-specific skills (Hegarty and Waller, 2005) such as interpreting map elements, reasoning over spatial relations, and integrating evidence across multiple maps, as well as broader capabilities emphasized in LVLM research that include text grounding (Singh et al., 2019; Sidorov et al., 2020), numeric and logical inference (Lu et al., 2024a; Hu et al., 2023), multi-image integration (Wang et al., 2024a; Xia et al., 2025), and retrieval (Wang et al., 2025a; 2024b). A cartographic reasoning benchmark can therefore probe comprehensive reasoning and provide a clear understanding of the spatial intelligence of LVLMs.
@@
A growing line of work began to evaluate LVLMs on map VQA, yet these benchmarks do not fully assess cartographic reasoning. Earlier datasets pose chart-style questions that can be answered without interpreting spatial relations, which bypasses the topological, metric, and directional inferences that are central to map comprehension (Koukouraki et al., 2025; Chang et al., 2022). Other efforts cover only a subset of relations as they target specific tasks such as navigation (Feng et al., 2025; Kazemi et al., 2025) or entity identification (Dihan et al., 2025). While suitable for those objectives, such coverage is insufficient for evaluating human-like map understanding (Liben et al., 2010). Many benchmarks also restrict the stylistic variability of maps. Some focuses on choropleths (Koukouraki et al., 2025; Chang et al., 2022; Mukhopadhyay et al., 2025), others rely on maps created with map-coloring tools (Srivastava et al., 2025) or common web basemaps (Kazemi et al., 2025; Dihan et al., 2025). Several others focus on limited thematic domains (e.g., geology (Huang et al., 2025)) or on restricted geographic coverage (Chang et al., 2022; Srivastava et al., 2025). These constraints overlook the heterogeneity in toponyms, labeling conventions, projections, and symbology that real-world cartography demands (Slocum et al., 2022; Robinson et al., 1978). Multi-map reasoning is rarely evaluated, with limited exceptions (Kazemi et al., 2025), even though practical cases often require integrating evidence across multiple maps (e.g., reconciling transit maps with future land-use maps for urban planning) and aligning overlapping information (Lupien and Moreland, 1987). Moreover, although document-level multimodal understanding is emphasized in other LVLM benchmarks, existing map VQA benchmarks seldom require selecting the correct map among many images in long reports, despite government documents and technical documents containing numerous, visually similar, context-dependent maps (Federal Emergency Management Agency, 2025; U.S. Environmental Protection Agency, 2025; SEDAR+, 2025). As a result, current map VQA settings underestimate the demands of comprehensive map understanding, leaving it unclear whether LVLMs possess human-like map-reading competencies. Full cartographic reasoning remains beyond the scope of what existing VQA benchmarks assess.
@@
We introduce FRIEDA, a benchmark designed for evaluating multi-map, multi-step, comprehensive cartographic reasoning in LVLMs. We curate maps from public documents of various thematic domains (e.g., geological surveys, planning reports, environmental studies) to develop questions that require models to interpret maps as they appear in reports, mirroring practical scenarios in which a reader must synthesize evidence from maps embedded in a document (see Figure 1). The collection encompasses a diverse range of styles, projections, and scales. We create each question such that it requires (1) reasoning over topological, metric, and directional relations, (2) interpreting map elements and their semantics (e.g., legends, map scales, and north arrows), and, when applicable, (3) integrating information across multiple maps, and (4) selecting the appropriate map(s) from a document to answer the query. To probe genuine reasoning rather than random guessing, the answers are in a free-form (not multiple-choice) format. The benchmark evaluation comprises two settings: a direct setting, in which the model is given the relevant map images along with the question to evaluate map comprehension, and a contextual setting, in which the model must first retrieve the correct maps from a broader within-document collection before answering. We score outputs using a unified, task-aware protocol aligned to the three spatial-relation categories. We evaluate textual responses (topological and semantic labels) with LLM-as-Judge (Gu et al., 2025), distance responses (numeric values with units) with unit-aware parsing and mean absolute percentage error (MAPE), and directional responses (cardinal directions for relative position) with angular tolerance over the eight directions. We compare the result against the human upper bound derived from multi-annotator agreement to contextualize LVLM performance. By aligning our tasks with the competencies expected of human map-readers (Goodchild, 2012; Liben et al., 2010) and explicitly targeting compositional cross-image inference that is largely absent from prior map VQA work, FRIEDA fills a crucial gap in state-of-the-art LVLM evaluation.
@@ 2 Task Definition
Spatial relations describe how geographic features relate in space (Carlson and Logan, 2001), how they are positioned in space (Majic et al., 2021), and how their geometries interact (Renzhong, 1998). In geographic information systems (GIS) and spatial cognition, these relations are often grouped into three categories: topological, metric, and directional (Cohn and Renz, 2007; Clementini et al., 1993). To make these abilities measurable and comparable, FRIEDA separates questions by spatial relation type and grounds the topological portion in the 9-intersection model (Clementini et al., 1993). We consolidate finer-grained subtypes into their broader categories (e.g., cross classified as intersect, and contain classified as within), yielding four topological classes: border (shared boundary between regions), equal (coincident geometries), intersect (crossing or overlap of features), and within (containment or inclusion of one area inside another). We complement these with one metric primitive, distance, and one directional primitive, orientation. Together, these six relations maintain the expressiveness of spatial queries while aligning with users’ intuitive spatial reasoning.
@@
Maps are symbolic representations that encode spatial information through abstract conventions (Slocum et al., 2022). Therefore, interpreting map elements is a distinct skill central to cartographic reasoning. The key elements we target are map text (place and feature names), legends (mappings from color, icons, and patterns to semantic classes), map scales (measurements that convert the map distance to the real-world distance), and the compass (ESRI, 2021). The styles of these components vary widely across maps: map texts may use different typography or placement rules (Monmonier, 2015), legends may use continuous color ramps or discrete pictograms (Slocum et al., 2022), map scales may appear as bars or frames around the map (Robinson, 1995), and the compass may be a compass rose or a north arrow (Slocum et al., 2022). Practical map interpretation requires grasping the concepts of map elements rather than simply recognizing their shapes. Consequently, our design includes questions that require reading map texts, decoding legends, using the map scale, and applying orientation to demonstrate true map literacy by linking abstract visual encodings to their underlying semantics.
@@
Beyond interpreting spatial relations and map elements, practitioners regularly perform cross-map comparison and fusion to synthesize multiple map editions or thematic layers (Lupien and Moreland, 1987). Our multi-map setting reflects this practice: we curate questions that present two or more maps together and require the model to integrate evidence by aligning shared symbols, reconciling differences in labels, map scales, and orientation, and identifying co-referent regions or features (Foody, 2007). Extracting distributions and patterns is widely recognized as a core capability (Ishikawa, 2016; Rexigel et al., 2024; Morita and Fukuya, 2025). By testing this setting, we move beyond isolated spatial computation to evaluate deeper cartographic reasoning across varied depictions of the same space.
@@
To mirror practical workflows (Mathew et al., 2021; Tanaka et al., 2023), we evaluate a contextual setting (FRIEDA-contextual) in which a model must identify the relevant map before answering a question. In this scenario, we provide the model with multiple maps from the same source (i.e., a document), and the model must perform within-document retrieval using cues in the maps, such as titles, legends, or labels. By evaluating model performance on FRIEDA-contextual, we capture a core aspect of real map use: the model must understand how map elements encode meaning and leverage that understanding to select the required map from thematically related alternatives that vary in data layers, geographic extent, or purpose (Ishikawa, 2016).
@@ 3.1 Benchmark Statistics
Table 1 shows that FRIEDA comprises 17,030 map images drawn from 210 documents. To capture real-world variability, these maps span diverse geographies (32 countries) and six thematic domains, exhibiting heterogeneous styles, including varied color palettes, legends, and symbol conventions. The benchmark comprises 500 questions, of which 202 are single-map and 298 are multi-map. The multi-map subset consists primarily of two-map comparisons (295 questions), with a few cases requiring reasoning across three (2 questions) or four (1 question) maps.
@@
For FRIEDA-contextual, we provide between 2 and 9 irrelevant maps in addition to the relevant map(s) required to answer the question. The total number of input images averages 9.54 maps ($\sigma=1.27$) per question across both the single- and multi-map settings, with an average relevant-to-irrelevant maps ratio of 1:5.71.
@@ 3.2 Benchmark Construction
To capture stylistic and geographic diversity, we curate maps from publicly available government and multilateral reports across domains including geology (SEDAR+, 2025), national park management (National Park Service, 2025), environmental assessments (U.S. Environmental Protection Agency, 2025; Environmental Protection Agency, 2025; Ministry of Agriculture Climate Change and Environment, 2025), disaster response (Federal Emergency Management Agency, 2025), urban planning (Seattle Planning and Community Development, 2024; Department of Community Development, 2025; Urban Redevelopment Authority, 2025; City of Cape Town, 2025), and infrastructure investment (AIIB, 2025). We limit the sources to documents written in Latin characters to focus on cartographic reasoning rather than translation. We extract images using Idefics3-8B (Laurençon et al., 2024) with a custom prompt (Appendix B.1) and manually verify that each extracted set contains only cartographic maps (examples of excluded non-maps appear in Appendix B.1.1). To support FRIEDA-contextual, we retain only documents with at least four maps. We select contextual maps based on their page proximity to the target map; this ensures they are thematically and stylistically related to the target maps. We then shuffle the map order to prevent LVLM from using positional cues to identify the target maps.
@@
For each collected map, we use GPT-4 and GPT-o3 (see Appendix B.6) to propose candidate questions, their corresponding targeted spatial relations, and reference answers. Rather than enforcing fixed templates, we accept any phrasing that unambiguously expresses the intended relation to reflect various forms of paraphrases (e.g., “Is A within B?” vs. “Does B contain A?”). To ensure the benchmark tests visual cartographic reasoning over memorization or web retrieval, we screen each question for searchability using GPT with web search enabled. In addition, we discard questions that can be answered without inspecting the map image. We label each question’s answer type as textual (short span or categorical label), distance (numeric with units), or direction to enable analysis by response modality.
@@
All LLM-proposed candidate questions undergo a pre-annotation curation stage. The two question curators (one with 7 years of GIS experience and another with 2 years of experience in geospatial data) manually verify gold answers against source maps and rewrite or discard ambiguous questions. This step ensures that FRIEDA consists solely of clear, high-quality questions prior to the broader validation phase with the annotator.
@@
We validate each question using annotations from 11 Ph.D. researchers (eight with map expertise), collected over four weeks. Annotators confirm that each question is answerable from the provided map(s) and, for multi-map questions, verify that all maps are required to answer the question. To prevent bias, curators do not validate their own edits. We retain the question only if the majority ($\geq$ 2/3) agrees with the gold answer. In a rare case (currently two questions in FRIEDA) where all three annotators agree on an answer that contradicts the gold answer, we conduct a secondary review to update the gold answer if consensus is reached. In total, we remove 61 questions that do not meet the agreement threshold of $\geq$ 2/3. Appendix B.3 details the instruction prompt provided to the annotators, and Appendix B.5 shows the annotation interface.